Write the equation of the line that passes through point (-5, -3) with a slope of 2.
Answer by checkley75(3666)

Y=mX+b IS THE LINE EQUATION WHERE X&Y ARE A SET OF POINTS & (m)=SLOPE & b=Y INTERCEPT.
-3=2*-5+b
-3=-10+b
b=-3+10
b=7 Y INTERCEPT. THUS THE LINE EQUATION IS:
Y=2X+7